OpenStudy (seattle12345):

Help please.... Solve the proportion: 9/(w-5)=2/11

OpenStudy (zzr0ck3r):

w-5 = 11/2 * 9 w = 11/2 * 9 +5

OpenStudy (zzr0ck3r):

w = (11/2) * 9 + 5

OpenStudy (seattle12345):

109/2 right?

OpenStudy (zzr0ck3r):

i think its 99/2 + 10/2

OpenStudy (lgbasallote):

lol yes you're right @seattle12345 109/2 = 99/2 + 10/2 ^_^

OpenStudy (zzr0ck3r):

yes, plug it in for w and check